# Rapid, strongly magnetized accretion in the zero-net-vertical-flux shearing box

### v1

@@ 6.4 The dynamo loop with net vertical flux
The dynamo mechanism with net vertical flux requires separate attention, both because of previous results and because it operates differently. In Fig. 15 we plot (for $\beta_{z}=100$), the dynamo contributions from Eq. 10 for both $\overline{B_{y}}$ (as in Fig. 12) and $\overline{B_{x}}$, each normalized as before to show the relative contribution per orbit at each $z$. The $\overline{B_{y}}$ balance shows significantly stronger outflows than in Fig. 12, balanced by a larger shear-generation term, $-q\Omega\overline{B_{x}}$ due to its strong $\overline{B_{x}}$ (this also causes the large mean magnetic field $\alpha_{\overline{B}}\approx\alpha$ contribution in Fig. 9f). Fluctuations, which have a similar amplitude to the ZNVF low-$\beta$ states, play little role in the balance (see Fig. 9e). In the lower panel, we see that $\overline{B_{x}}$ is also being sourced in a laminar way, via stretching of the mean $\beta_{z}=100$ vertical field by a self-induced vertically sheared radial flow ($\overline{B_{z}}\partial_{z}\overline{U_{x}}$).
@@
This behavior, which is clearly different from that at ZNVF, is simply caused by a single, box-scale MRI channel mode developing across the domain. Indeed the relative signs, symmetries, and sizes of $\overline{\bm{B}}$ and $\overline{\bm{U}}$ in this $\beta_{z}=100$ state are similar to those of a growing MRI mode. The solution is similar to that discussed in the contexts of winds in Lesur et al. (2013) but with the opposite symmetry: here and in Bai & Stone (2013), $\overline{B_{y}}$ has a peak in the midplane (like $\cos(z)$); for the solutions in Lesur et al. (2013), $\overline{B_{y}}$ passes through zero at $z=0$ (like $\sin(z)$). By extension, effectively all of the angular momentum stress comes from this channel mode, explaining to the strong dominance of $\alpha_{\overline{B}}$ compared to the turbulent contributions.
@@
An additional important feature is that the symmetry of this $\overline{U_{x}}$ may not be possible to realize in global simulations. In particular, as pointed out by Bai & Stone (2013) in the context of the outflow/wind (see their figure 12), this $\partial_{z}\overline{U_{x}}>0$ requires that $\overline{U_{x}}$ and $\overline{B_{x}}/\overline{B_{z}}$ have opposite signs above and below the midplane, implying a radial inflow for $z>0$ and outflow for $z<0$ or vice versa (depending on the sign of $B_{z}$). The flow is also fast ($\overline{U_{x}}\gtrsim c_{s}$ at larger $z$; Bai & Stone, 2013) thereby requiring a strongly asymmetrical corona that seems at odds with those expected and observed in corresponding global simulations. Indeed, rather than the single-signed strong toroidal field seen here and in past local studies, Zhu & Stone (2018); Mishra et al. (2020) observe a sign flip in $\overline{B_{y}}$ (and $\overline{B_{x}}$) across the midplane, with approximately symmetric inflow both above and below the midplane. We note that this state still appears to have the structure of a single large-scale channel mode, albeit one with the opposite symmetry like those studied by Lesur et al. (2013). This change causes $\beta\gtrsim 1$ in the midplane of the global solution of Mishra et al. (2020) at $\beta_{z}=100$, although the accretion stress above and below remains mean-field dominated. While further study of these solutions is clearly needed, the analysis helps to elucidate the cause of the strong influence of net flux on both local and global simulations — it destabilizes a box-scale MRI mode, which then dominates the generation of the mean fields and angular momentum transport, driving strong outflows.
@@ 7 Discussion & Conclusions
In this work, we have introduced and characterized a novel strongly magnetized state of self-sustaining accretion disk turbulence in the local stratified shearing box. The system maintains a strong mean azimuthal field in the midplane, with $\beta\lesssim 1$, modestly sub-Alfvénic fluctuations ($|\delta\bm{B}|\lesssim\overline{B_{y}}$), and large accretion stresses, $\alpha\simeq 0.5$. A net vertical flux — traditionally thought to be necessary to sustain such large $\alpha$ — is not required, although the field structure and turbulence properties do superficially resemble previous NVF results (e.g., Bai & Stone, 2013) even though the physics of how the ZNVF is sustained is rather different. So long as the vertical domain size is sufficiently large (in units of $H_{\rm th}$), the state is also robust to choices of vertical boundary conditions and numerical options — indeed, we show explicitly that the azimuthal flux is continuously escaping through the boundaries. Our results resemble those of Johansen & Levin (2008) with the crucial difference that in their case the total flux is constrained by the periodic boundary conditions.
@@
We argue that the transition to this low-$\beta$ state, as opposed to the well-known high-$\beta$ cyclic dynamo discussed in many previous works (e.g., Davis et al., 2010; Simon et al., 2012), is controlled by the initial magnetization of the midplane: if there is sufficient azimuthal flux after the initial conditions rearrange for a $\beta\lesssim 1$ midplane and $\beta\ll 1$ corona, the system can maintain the low-$\beta$ state; if there is not, it inevitably loses all midplane flux, transitioning into high-$\beta$ dynamo cycles. The initial azimuthal (or radial) flux thus acts like a “switch” between two states of vastly different $\alpha$ in a system with otherwise identical parameters. Empirically we find that an initial $\beta\simeq 0.1$ in the initial Gaussian toroidal field is the dividing line between the disk evolving to the low-$\beta$ or high-$\beta$ state. Given the well-known convergence and microphysics dependence of ZNVF shearing box results in the high-$\beta$ state (Fromang et al., 2007; Ryan et al., 2017), the transition between the two states found here, and indeed the properties and existence of the low-$\beta$ state, may also be sensitive to these physical and numerical parameters. This should clearly be studied more in future work, where mesh refinement may be helpful in allowing the thermal midplane to be better resolved than has been possible here. The importance of turbulent magnetic flux transport in determining the properties of the low-$\beta$ state also specifically motivates future simulations with explicit resistivity and varying magnetic Prandtl number.
@@
Given the variety of rapid state transitions observed in X-ray transients, dwarf novae (e.g. Smak, 1984) and changing-look AGN (Ricci & Trakhtenbrot, 2023), the strong dependence of the efficiency of angular momentm transport on the “initial” magnetic field may be of particular interest. For instance, in dwarf novae, $\alpha$ is seen to vary in between $\alpha\sim 0.1$-$0.4$ in the hot state and $\alpha\sim 0.01$-$0.03$ in the cold state (King et al., 2007; Kotko & Lasota, 2012), rather similar to the difference seen here between the low- and high-$\beta$ states, while X-ray binaries may exhibit yet higher $\alpha\sim 0.2$-$1$ (Tetarenko et al., 2018). While changes in the net vertical flux, thermal and ionization states, and/or winds can potentially explain this variability (e.g., Gammie & Menou, 1998; King & Ritter, 1998; Hirose et al., 2014; Begelman et al., 2015; Scepi et al., 2018; Tetarenko et al., 2018; Zhu & Stone, 2018), the possibility hinted at by our results has the nice feature that extra magnetic flux in almost any configuration (toroidal, radial, or poloidal) could be sufficient to cause significant changes in the efficiency of angular momentum transport. For instance, in feeding such a disk, if a parcel of gas with higher-than average magnetic field caused the toroidal field strength to cross the $\beta\lesssim 1$ threshold, it would trigger the transition from the high- to low-$\beta$ state. The same could occur if the disk undergoes a thermal collapse, amplifying the toroidal field by flux freezing. This would result in significantly more rapid accretion, potentially then depleting the disk of magnetic flux and density, resulting in an eventual transition back to the high-$\beta$ state. Of course, such a scenario is not necessarily at odds with previous ideas; its benefit is to add a clear bimodality to the system, causing large, sudden changes to $\alpha$ due to small changes in disk parameters.
@@ 7.1 The cause of the state transition
In the course of analysing the low-$\beta$ state and its transition to the high-$\beta$ state, we have run a wide variety of simulations with different initial conditions and other options, leading to various more detailed conclusions. These are listed in §2, to which we refer the reader for a more in depth summary of our computational results. Forgoing the summary here, we instead attempt to propose a scenario for the key physical ingredients that create bi-modality, causing the appearance of the two distinct (low- and high-$\beta$) states with no intermediate state. This remains far from certain — indeed our attempts at diagnosing the field-generation mechanism remained partially unsuccessful (see §6.3) — but may provide a helpful framework for further development. The scenario has three main elements:
@@
The disk midplane and its atmosphere are highly unstable, exhibiting strong trans- and super-sonic turbulence. This rapidly transports flux upwards, implying that in order to sustain itself, the system must continuously generate flux through a (non-cyclic) large-scale dynamo. The flux is regenerated and removed on several-orbit timescales, as shown in Figs. 12 and 13, implying the dynamo must be efficient and robust. Given that the mean shear naturally generates strong azimuthal flux from radial flux, the turbulent fluctuations must generate radial flux from azimuthal flux to “close the loop.”
@@
Both the Parker instability and MRI seem to play important roles in maintaining the low-$\beta$ state, as evidenced by our simple turbulence model (Fig. 6) and previous works (Johansen & Levin, 2008). However, evidence suggests that fluctuations driven by the Parker instability generically and naturally generate azimuthal from radial flux, with MRI playing a subsidiary role. First, even when the system is not in steady state with $\beta$ increasing in time, the dynamo still operates (flux is just being lost faster than it is created; Fig. 13); in this regime $\beta\ll 1$ so the MRI is suppressed (likewise $\beta\ll 1$ resolution-limited states sustain a dynamo). Second, in the midplane ($|z|\lesssim H_{\rm th}$) the net radial field is absent or of the wrong sign to support the dynamo (Figs. 1 and 12); this is presumably where MRI drives turbulence, since the Parker instability is effectively stable. Together these suggest that Parker-driven fluctuations drive the dynamo itself, while MRI is likely needed to drive midplane turbulence that can diffuse $\overline{B_{y}}$ to maintain the strongly magnetized midplane.
@@
Together with points (1) and (2) — that the Parker instability is needed for sustaining the azimuthal flux, which rapidly escapes if it is not continuously regenerated — point (3) implies that the Parker-induced feedback would become less efficient at regenerating the mean field once $\beta_{\rm mid}\gtrsim 1$. This would then further raise $\beta$ by its inability to regenerate more flux, running away until the flux is completely expelled from the midplane. Thus results the high-$\beta$ state, which presumably sustains turbulence via a different midplane mechanism (unrelated to the Parker instability) exhibiting characteristic ${\sim}10$-orbit reversal of $\overline{\bm{B}}$, $\delta\bm{B}\gg\overline{\bm{B}}$, and $P_{B}\ll P_{\rm th}$. This scenario thus qualitatively explains why the system switches states based on the initial flux, with no “intermediate” state available (see Fig. 4) — as $\beta_{\rm mid}$ rises above unity, the feedback mechanism becomes continuously weaker. Further stipulating that role of the MRI is primarily to provide a source of midplane turbulence, the scenario is also consistent with the behavior at $\beta\ll 1$, in particular (i) the inability to maintain a $\beta\ll 1$ steady state despite the similar dynamo, since the MRI cannot maintain supersonic turbulence at $z=0$ (where Parker shuts off), as would be needed to maintain a density scale height ${\gg}H_{\rm th}$; and (in contrast) (ii) the ability to maintain a $\beta\ll 1$ in steady state if limited by resolution — with a poorly resolved midplane, the physical (turbulent) diffusion is replaced by numerical diffusion, which stops the collapse of the density.
@@ 7.2 Subsidiary results
In addition to the main results of this work, discussed above, we have noted various other subsidiary results relating to the high-$\beta$ dynamo cycles and accreting states with net vertical flux:
@@
Periodic reversals of $\overline{B_{y}}$, which have been traditionally associated with the MRI, appear to persist in the absence of rotation, albeit with some morphological differences (see Fig. 8). This does not seem to disagree with previous works, which mostly have not explored this particular numerical experiment of setting $\Omega=0$, but may be of interest for understanding the mechanism via which this turbulence self sustains (in idealized dynamo-focused simulations, cyclic behavior is also observed without rotation, so this is perhaps not unexpected; e.g., Brandenburg et al. 2008; Squire & Bhattacharjee 2015b). Similar cycles are also observed with a net vertical flux in the absence of rotation.
@@
With strong net vertical flux ($\beta_{z}\lesssim 100$), the large angular momentum transport is almost entirely driven by the mean magnetic-field stress ($\alpha_{\overline{B}}=-\overline{B_{x}}\,\overline{B_{y}}$; Bai & Stone, 2013), with $\overline{B_{x}}$ and $\overline{B_{y}}$ maintained by the mean flows ($\partial_{z}\overline{U_{x}}$ stretches $\overline{B_{z}}$ into $\overline{B_{x}}$; the Keplerian flow stretches $\overline{B_{x}}$ into $\overline{B_{y}}$). This structure is effectively a single MRI channel mode stretching across the full vertical extent of the domain, although its appearance is rather similar to the low-$\beta$ ZNVF state. As also noted by Bai & Stone (2013) in the context of the winds produced by their simulations, the symmetry of the solution likely implies it cannot occur in more realistic, global settings, since it requires strong (supersonic) oppositely directed flows above and below the midplane in order to sustain its strong mean fields and thus its large $\alpha$. Indeed, unlike the shearing box, which exhibits a single-signed $\overline{B_{y}}$ at any particular time, similar ($\beta_{z}\approx 100$) global NVF simulations develop strong ($\beta\gtrsim 1$) mean fields that reverse their polarity across the midplane (Mishra et al., 2020; see also Zhu & Stone 2018). We note in passing that this global structure in Zhu & Stone (2018); Mishra et al. (2020) is also essentially a single channel mode, but with the opposite symmetry about $z$: effectively $\overline{B_{y}}\propto\sin(z)$ and $\overline{U_{x}}\propto\cos(z)$ in the global domain, rather than $\overline{B_{y}}\propto\cos(z)$ and $\overline{U_{x}}\propto-\sin(z)$ in the local one. This change allows the global system to maintain a symmetric flow structure across the midplane.
@@
7.3 Application to global models

### v3

@@ 6.4 The dynamo loop with net vertical flux
The dynamo mechanism with net vertical flux requires separate attention, both because of previous results and because it operates differently. In Fig. 15 we plot (for $\beta_{z}=100$), the dynamo contributions from Eq. 10 for both $\overline{B_{y}}$ (as in Fig. 12) and $\overline{B_{x}}$, each normalized as before to show the relative contribution per orbit at each $z$. The $\overline{B_{y}}$ balance shows significantly stronger outflows than in Fig. 12, balanced by a larger shear-generation term, $-q\Omega\overline{B_{x}}$ due to its strong $\overline{B_{x}}$ (this also causes the large mean magnetic field $\alpha_{\overline{B}}\approx\alpha$ contribution in Fig. 9e). Fluctuations, which have a similar amplitude to the ZNVF low-$\beta$ states, play little role in the balance (see Fig. 9d). In the lower panel, we see that $\overline{B_{x}}$ is also being sourced in a laminar way, via stretching of the mean $\beta_{z}=100$ vertical field by a self-induced vertically sheared radial flow ($\overline{B_{z}}\partial_{z}\overline{U_{x}}$).
@@
This behavior, which is clearly different from that at ZNVF, is simply the manifestation of a single, box-scale MRI channel mode developing across the domain. Indeed the relative signs, symmetries, and sizes of $\overline{\bm{B}}$ and $\overline{\bm{U}}$ in this $\beta_{z}=100$ state are similar to those of a growing MRI mode. The solution resembles that discussed in the context of winds in Lesur et al. (2013) but with the opposite symmetry: here and in Bai & Stone (2013), $\overline{B_{y}}$ has a peak in the midplane (like $\cos(z)$); for the solutions in Lesur et al. (2013), $\overline{B_{y}}$ passes through zero at $z=0$ (like $\sin(z)$). By extension, effectively all of the angular momentum stress comes from this channel mode, explaining the strong dominance of $\alpha_{\overline{B}}$ compared to the turbulent contributions.
@@
An additional important feature is that the symmetry of this $\overline{U_{x}}$ may not be possible to realize in global simulations. In particular, as pointed out by Bai & Stone (2013) in the context of the outflow/wind (see their figure 12), this $\partial_{z}\overline{U_{x}}>0$ requires that $\overline{U_{x}}$ and $\overline{B_{x}}/\overline{B_{z}}$ have opposite signs above and below the midplane, implying a radial inflow for $z>0$ and outflow for $z<0$ or vice versa (depending on the sign of $B_{z}$). The flow is also fast ($\overline{U_{x}}\gtrsim c_{s}$ at larger $z$; Bai & Stone, 2013) thereby requiring a strongly asymmetrical corona that seems at odds with those expected and observed in corresponding global simulations. Indeed, rather than the single-signed strong toroidal field seen here and in past local studies, Zhu & Stone (2018); Mishra et al. (2020) observe a sign flip in $\overline{B_{y}}$ (and $\overline{B_{x}}$) across the midplane, with approximately symmetric inflow both above and below the midplane. We note that this state still appears to have the structure of a single large-scale channel mode, albeit one with the same symmetry as those studied by Lesur et al. (2013). This change causes $\beta\gtrsim 1$ in the midplane of the global solution of Mishra et al. (2020) at $\beta_{z}=100$, although the accretion stress above and below remains mean-field dominated. While further study of these solutions is clearly needed, the analysis helps to elucidate the cause of the strong influence of net flux on both local and global simulations — it destabilizes a box-scale MRI mode, which then dominates the generation of the mean fields and angular momentum transport, driving strong outflows.
@@ 7 Discussion & Conclusions
In this work, we have introduced and characterized a novel strongly magnetized state of self-sustaining accretion disk turbulence in the local stratified shearing box. The system maintains a strong mean azimuthal field in the midplane, with $\beta\lesssim 1$, modestly sub-Alfvénic fluctuations ($|\delta\bm{B}|\lesssim\overline{B_{y}}$), and large accretion stresses, $\alpha\simeq 0.5$. A net vertical flux — traditionally thought to be necessary to sustain such large $\alpha$ — is not required, although the field structure and turbulence properties do superficially resemble previous NVF results (e.g., Bai & Stone, 2013) even though the sustenance mechanism of the ZNVF state is different. So long as the vertical domain size is sufficiently large (in units of $H_{\rm th}$), the state is also robust to choices of vertical boundary conditions and numerical options — indeed, we show explicitly that the azimuthal flux is continuously escaping through the boundaries.
@@
We argue that the transition to this low-$\beta$ state, as opposed to the well-known high-$\beta$ cyclic dynamo discussed in many previous works (e.g., Davis et al., 2010; Simon et al., 2012), is controlled by the initial magnetization of the midplane: if there is sufficient azimuthal flux after the initial conditions rearrange to create a $\beta\lesssim 1$ midplane and $\beta\ll 1$ corona, the system can maintain the low-$\beta$ state; if there is not, it inevitably loses all midplane flux, transitioning into high-$\beta$ dynamo cycles. The initial azimuthal (or radial) flux thus acts like a “switch” between two states of vastly different $\alpha$ in a system with otherwise identical parameters. Empirically we find that an initial $\beta\simeq 0.1$ in the initial Gaussian toroidal field is the dividing line between the disk evolving to the low-$\beta$ or high-$\beta$ state. Given the well-known convergence and microphysics dependence of ZNVF shearing box results in the high-$\beta$ state (Fromang et al., 2007; Ryan et al., 2017), the transition between the two states found here, and indeed the properties and existence of the low-$\beta$ state, may also be sensitive to these physical and numerical parameters. This should clearly be studied more in future work, where mesh refinement may be helpful in allowing the thermal midplane to be better resolved than has been possible here. The importance of turbulent magnetic flux transport in determining the properties of the low-$\beta$ state also specifically motivates future simulations with explicit resistivity and varying magnetic Prandtl number.
@@
Given the variety of rapid state transitions observed in X-ray transients, dwarf novae (e.g. Smak, 1984) and changing-look AGN (Ricci & Trakhtenbrot, 2023), the strong dependence of the efficiency of angular momentum transport on the “initial” magnetic field may be of particular interest. For instance, in dwarf novae, $\alpha$ is seen to vary in between $\alpha\sim 0.1$-$0.4$ in the hot state and $\alpha\sim 0.01$-$0.03$ in the cold state (King et al., 2007; Kotko & Lasota, 2012), rather similar to the difference seen here between the low- and high-$\beta$ states, while X-ray binaries may exhibit yet higher $\alpha\sim 0.2$-$1$ (Tetarenko et al., 2018). While changes in the net vertical flux, the MRI, and/or winds can potentially explain this variability (e.g., Gammie & Menou, 1998; Hirose et al., 2014; Begelman & Armitage, 2014; Scepi et al., 2018; Zhu & Stone, 2018), as well as mechanisms unrelated to changes in the magnetic field (e.g., King & Ritter, 1998; Scepi et al., 2024), the possibility hinted at by our results has the nice feature that extra magnetic flux in almost any configuration (toroidal, radial, or vertical) could be sufficient to cause significant changes in the efficiency of angular momentum transport. For instance, in feeding such a disk, if a parcel of gas with higher-than average magnetic field caused the toroidal field strength to cross the $\beta\lesssim 1$ threshold, it would trigger the transition from the high- to low-$\beta$ state. As seen in Machida et al. (2006), whose global simulations reach a $\beta\sim 0.1$ state with similar properties to that studied here, the same could occur if the disk undergoes a thermal collapse, amplifying the toroidal field by flux freezing. This would result in significantly more rapid accretion, potentially then depleting the disk of magnetic flux and density to cause an eventual transition back to the high-$\beta$ state. This scenario is not necessarily at odds with various previous proposals; the new insight here is that the magnetic-field strength adds bimodality to the local turbulence, with large, sudden changes to $\alpha$ resulting from small changes in disk parameters. However, further study with a more detailed treatment of radiation is certainly needed, for instance to understand whether such transitions, which occur over tens of orbital periods in our simulations, disagree with observed smoother transitions through intermediate states (e.g., Skipper & McHardy, 2016).
@@ 7.1 The cause of the state transition
In the course of analysing the low-$\beta$ state and its transition to the high-$\beta$ state, we have run a wide variety of simulations with different initial conditions and other options, leading to various more detailed conclusions. These are listed in §2, to which we refer the reader for a more in depth summary of our computational results. Forgoing a summary here, we instead propose a scenario for how the system attains bimodality, giving rise to the distinct low- and high-$\beta$ states. This physics remains far from certain — indeed our attempts at diagnosing the field-generation mechanism remained partially unsuccessful (see §6.3) — but may provide a helpful framework for further development. The scenario has three main elements:
@@
The disk midplane and its atmosphere are highly unstable, exhibiting strong trans- and super-sonic turbulence. This rapidly transports flux upwards, implying that in order to sustain itself, the system must continuously generate flux through a (non-cyclic) large-scale dynamo. The flux is regenerated and removed on several-orbit timescales, as shown in Figs. 12 and 13, implying the dynamo must be efficient and robust. Given that the mean shear generates strong azimuthal flux from radial flux, the turbulent fluctuations must generate radial flux from azimuthal flux to “close the loop.”
@@
Both the Parker instability and MRI seem to play important roles in maintaining the low-$\beta$ state, as evidenced by the similarity of the midplane turbulence to the high-$\beta$ state (Fig. 6), our simple turbulence model (Fig. 7), and previous works (Johansen & Levin, 2008; Held et al., 2024). However, evidence suggests that fluctuations driven by the Parker instability generically and naturally generate azimuthal from radial flux, with MRI playing a subsidiary role. First, even when the system is not in steady state with $\beta$ increasing in time, the dynamo still operates (flux is just being lost faster than it is created; Fig. 13); in this regime $\beta\ll 1$ so the MRI is suppressed (likewise $\beta\ll 1$ resolution-limited states sustain a dynamo). Second, in the midplane ($|z|\lesssim H_{\rm th}$) the net radial field is absent or of the wrong sign to support the dynamo (Figs. 2 and 12); this is presumably where MRI drives turbulence, since the Parker instability is stable. Together these suggest that Parker-driven fluctuations drive the dynamo itself, while MRI is likely needed to drive midplane turbulence that can diffuse $\overline{B_{y}}$ to maintain the strongly magnetized midplane.
@@
Together with points (1) and (2) — that the Parker instability is needed for sustaining the azimuthal flux, which rapidly escapes if it is not continuously regenerated — point (3) implies that the Parker-induced feedback becomes less efficient at regenerating the mean field for $\beta_{\rm mid}\gtrsim 1$. This would then further raise $\beta$ by its inability to regenerate more flux, running away until the flux is completely expelled from the midplane. Thus results the high-$\beta$ state, which presumably sustains turbulence via a different midplane mechanism (unrelated to the Parker instability) exhibiting characteristic ${\sim}10$-orbit reversal of $\overline{\bm{B}}$, $\delta\bm{B}\gg\overline{\bm{B}}$, and $P_{B}\ll P_{\rm th}$. This scenario thus qualitatively explains why the system switches states based on the initial flux, with no “intermediate” state available (see Fig. 5) — as $\beta_{\rm mid}$ rises above unity, the feedback mechanism becomes continuously weaker. If we further assume that the primary role of the MRI is to generate midplane turbulence, the scenario remains consistent with the behavior at $\beta\ll 1$. Specifically: (i) a steady state with $\beta\ll 1$ cannot be maintained despite a similar dynamo, because the MRI fails to sustain supersonic turbulence at $z=0$ (where Parker instability shuts off), which would be required to support a density scale height $\gg H_{\text{th}}$; and (ii) in contrast, a steady state with $\beta\ll 1$ can persist if resolution is insufficient — when the midplane is poorly resolved, physical (turbulent) diffusion is replaced by numerical diffusion, preventing the collapse of the density.
@@ 7.2 Subsidiary results
In addition to the main results of this work, we have noted subsidiary results relating to the high-$\beta$ dynamo cycles and turbulence with net vertical flux:
@@
Periodic reversals of $\overline{B_{y}}$, which have been traditionally associated with the MRI, appear to persist in the absence of rotation, albeit with some morphological differences (see App. D, Fig. 19). This does not seem to disagree with previous works, which mostly have not explored the effect of setting $\Omega=0$, but may be of interest for understanding the mechanism via which this turbulence self sustains (idealized dynamo-focused simulations exhibit cyclic behavior without rotation, so this is perhaps not unexpected; e.g., Brandenburg et al. 2008; Squire & Bhattacharjee 2015b). Similar cycles are also observed with a net vertical flux in the absence of rotation (Fig. 20).
@@
With strong NVF ($\beta_{z}\lesssim 100$), the large angular momentum transport is almost entirely driven by the mean magnetic-field stress ($\alpha_{\overline{B}}=-\overline{B_{x}}\,\overline{B_{y}}$; Bai & Stone, 2013), with $\overline{B_{x}}$ and $\overline{B_{y}}$ maintained by the mean flows ($\partial_{z}\overline{U_{x}}$ stretches $\overline{B_{z}}$ into $\overline{B_{x}}$; the Keplerian flow stretches $\overline{B_{x}}$ into $\overline{B_{y}}$). This structure is effectively a single MRI channel mode stretching across the full vertical extent of the domain, although its appearance is rather similar to the low-$\beta$ ZNVF state. As also noted by Bai & Stone (2013) in the context of the winds produced by their simulations, the symmetry of the solution likely implies it cannot occur in more realistic, global settings, since it requires strong (supersonic) oppositely directed radial flows on either side of the midplane to sustain the mean fields and large $\alpha$. Indeed, unlike the shearing box, which exhibits a single-signed $\overline{B_{y}}$ at any particular time, similar ($\beta_{z}\approx 100$) global NVF simulations develop strong ($\beta\lesssim 1$) mean fields that reverse their polarity across the midplane (Mishra et al., 2020; see also Zhu & Stone 2018). We note in passing that this global structure in Zhu & Stone (2018); Mishra et al. (2020) is also essentially a single channel mode, but with the opposite symmetry about $z$: effectively $\overline{B_{y}}\propto\sin(z)$ and $\overline{U_{x}}\propto\cos(z)$ in the global domain, rather than $\overline{B_{y}}\propto\cos(z)$ and $\overline{U_{x}}\propto-\sin(z)$ in the local one. This change allows the global system to maintain a symmetric flow structure across the midplane.
@@
7.3 Application to global models and $\beta\ll 1$ fields